Evaluate:
a) 3^-2 - 3^0+(1/3)^-2
b)(3^-1+2^-1)^-2

Answer by stanbon(75887)

Evaluate:
a) 3^-2 - 3^0+(1/3)^-2
= (1/9) - 1 + 9
= 8 1/9
= 73/9
-------------------
b)(3^-1+2^-1)^-2
= ((1/3) + (1/2))^-2
= (5/6)^-2
= (25/36)^-1
= 36/25
= 1 11/25
===============
Cheers,
Stan H.